OpenStudy (maheshmeghwal9):

The no. of solutions of z^2+3z'=0 Are:-------? -->>Where, z=complex no. & z'=z's conjugate

OpenStudy (maheshmeghwal9):

How can we find the no. of solutions?Pleaz tell:)

OpenStudy (anonymous):

it has to do with the number of roots (the highest power). for this it is 2. z=-3,0

OpenStudy (maheshmeghwal9):

but how can i find it ,I mean how can I convert z' into z

OpenStudy (maheshmeghwal9):

and the answer is 4

OpenStudy (anonymous):

z=a+bi z'=a-bi

OpenStudy (maheshmeghwal9):

k! but the answer is 4.Why?

OpenStudy (anonymous):

a^2+b^2+3a-3bi=0 I didn't consider there are two squared unknowns for the complex square. That gives two values available for each of a and b.

OpenStudy (maheshmeghwal9):

k! thanx but there is only one equation in 2 variables , where is the second one ,coz we have to find the solutions

OpenStudy (maheshmeghwal9):

np:) if question is limited to finding the no. of solutions since the question is not asking the solutions it is only asking the no.s. but if we were said to find the solns then what would we do?

OpenStudy (anonymous):

I would solve by setting real and equal parts equal to their respective right side constants. a^2+b^2+3a=0 -bi=0i with nonzero values on the right, I would hope that with a and b both being quadratic, they would each have two roots.

OpenStudy (maheshmeghwal9):

k! thanx I got it .Are u doing engineering

OpenStudy (anonymous):

physics.

OpenStudy (maheshmeghwal9):

oh i see but from where?

OpenStudy (maheshmeghwal9):

I mean from where u are?

OpenStudy (anonymous):

Omaha, Nebraska

OpenStudy (maheshmeghwal9):

k! good luck :)

OpenStudy (anonymous):

U too

OpenStudy (maheshmeghwal9):

thanx!